# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

### v2

@@ 2 Technical Warm-Ups
Honeypot. Another common DEX scam is Honeypot [25]. The scammer maliciously designs the token’s smart contract to allow users to easily purchase the token but block them from selling or withdrawing it. The contract may include restrictions such as preventing sell transactions, allowing withdrawals only for whitelisted addresses, or imposing exorbitant transaction fees that make it impossible to retrieve funds. The scammer then exploits the liquidity pool through sell and withdraw actions, similar to the mechanics of a rug pull scam.
We defer more related work in Appendix A.
All the stated fraudulent behaviors share common characteristics regarding the liquidity pool: (i) the liquidity pool tokens (LPTs), which represent the initial liquidity [36], are not burned, and (ii) the ownership of the paired token remains entirely monopolized by the pool creator. Without implementing an ethical safeguard (i.e., burning initial LPTs), liquidity providers, including those involved in rug pulls, honeypots, and SLID scams (as identified in our findings), confront no financial risk (proven in Appendix B) for engaging in fraudulent activities. This enables them to directly drain assets from users’ investments, perpetuating the cycle of exploitation.
3 Motivation Example
This section discusses the behavior of Token $\hat{A}$ (all the collected data, including token names, addresses, and liquidity pool details, are anonymized for privacy & ethical reasons) with token address “0x4e84…a17c8”, created in August 2021, which serves as a motivational example for our study. This token was launched with a massive marketing campaign highlighting its high potential for generating profits for investors.
We have identified the liquidity pool of $\hat{A}$ as a SLID scam due to the deceptive actions of its owners. They slowly exploited the liquidity pool through a mixed sequence of token buying and selling on Uniswap while maintaining a stable price and trading volume during the initial months after its launch. The gradual profit-taking persisted for an extended period before the token owner executed a significant rug pull a year later. We briefly describe the sequence of actions taken by the owners in this cryptocurrency’s SLID scam.
① Attracting potential victims. To rapidly attract users, $\hat{A}$’s advertising campaign targeted buyers by promising huge token profits if purchased early. The owner engaged the community by promoting the token via their X account (formerly Twitter). They attracted 78.5K followers and collaborated with 20+ well-known crypto influencers. The campaign successfully drew 5,326 investors to purchase $\hat{A}$ on Uniswap within the first 70 days from the token’s LP deployment.
② Maintaining the pool and slowly taking profit.  The token’s liquidity pool was deployed on August 4th, 2021, with an initial owner liquidity deposit of \$19K. However, the liquidity pool for $\hat{A}$ was deployed without burning the owner’s initial tokens. This creates a persistent risk, as the owner of $\hat{A}$ can drain the liquidity pool and seize all funds invested by users to purchase $\hat{A}$ at any time. This is precisely what occurs in rug pull and honeypot scenarios (as described in §2).
We observed $\hat{A}$’s owner DEX activities (cf. Figure 2) over 70 days after deployment, including price and trading volume. During this period, the owner was engaged in a sequence of mixed activities, including buying and selling. Although these actions seem common and legitimate, our evaluation reveals that the selling amount consistently exceeded the buying amount (in paired tokens). Specifically, the total selling volume was nearly ten times larger, with the owner purchasing only $2.38\times 10^{15}$ tokens while selling $3.21\times 10^{16}$ tokens of $\hat{A}$ in total. This indicates that the owner possesses a significant number of tokens and attempts to incorporate this inflated amount into their selling orders. $\hat{A}$’s owner executed 139 buy and 136 sell actions, but with a nearly double volume gap that was worth \$783K and \$1.4M, respectively. This resulted in a huge profit for the pool owner of \$707K — nearly double the initial investment (of the base token amount the owner used to initially deposit and buy the paired token multiple times) — that comes directly from the investor’s money.
In contrast, in this 70-day period, those 5,326 investors committed with 10,374 buys, 3,898 deposits, 7,322 sells, but only 2 withdraw activities, with a total inflow of nearly \$1.35M (in base token) to the pool. The mix of buy and sell orders created the appearance of legitimacy in the token’s transaction history, helping to avoid the token price crash that would occur if they sold all at once (i.e., 1-day rug pull [37]).
③ Repeating the scam. While maintaining the appearance of token legitimacy, $\hat{A}$’s owner continued to sell only a small portion of the tokens compared to the overall LP value. The smallest and largest sell order values account for only 1% and 16% of the pool’s value. By carefully mixing buy and sell orders and controlling the selling amounts, the built-in scam detectors on DEXs are unable to identify this token as a scam. As of July 13, 2024, the scammer still maintains \$91K worth of liquidity in the pool and has remained inactive since their last recorded sell order on July 31, 2022.
Thus, despite the stable behavior of the token, a series of buying and selling actions (with smaller amounts) occurred at a slow pace, gradually adding an inflated amount of tokens to each sell order from the pool owner’s account. The sequence of actions resembles the profit-taking sell actions seen in rug pull, where an inflated amount is sold to drain the base tokens. This scam strategy can easily avoid investor attention while enabling the token to remain active and attract additional investments over an extended period.
4 Threat Model & Security Goal
Entities in DEXs. A typical DEX involves two entities: the pool owner (short for the owner), represented by the address that deploys the liquidity pool, and the investors/users, who participate in the pool’s activities.
Threat model. An SLID adversary (i.e., the pool owner) may exploit liquidity pools containing user/investor funds in DEXs by employing a gradual profit-taking strategy to: (i) evade the current fraud detection systems; and (ii) maintain user confidence, enabling larger profit over time. An adversary is supposed to have permissions for the following behaviors:
  • $\bullet$
    The pool owner can unethically monopolize the paired token by retaining their initial tokens instead of burning them when creating the liquidity pool. The owner can then extract profits directly from users’ investments/funds (in base tokens) and use profits to buy additional paired tokens.
  • $\bullet$
    The pool owner can perform activities including selling/withdrawing and buying/depositing at a relatively high frequency, each time involving a small amount of tokens. The price and volume for both tokens will not exhibit obvious fraudulent patterns. The owner can extract small portions of profit through these paired activities.
  • $\bullet$
    The pool owner commits to a long-term engagement in such activities, gradually draining funds while maintaining the pool’s appearance to avoid raising suspicion.
Security goals. We have the following three goals:
  • $\bullet$
    identify common features observed in SLID scams by analyzing sufficient real-world public data collected on Ethereum, addressing RQ1;
  • $\bullet$
    design a heuristic capable of detecting SLID fraudulent behavior from the liquidity pool’s visible activities (e.g., owner activity, profit-taking) and capture the occurred SLID fraud among Ethereum DEXs, addressing RQ2;
  • $\bullet$
    enhance detection robustness by enabling real-time identification of SLID scams. Building on the heuristic results, apply ML models to detect SLID scams during the early stages of liquidity pool operations, addressing RQ3.
5 Common Features of SLID Scam (RQ1)
In this section, we conduct an in-depth investigation into SLID liquidity pools. SLID scam is for the first time identified, where no detection methods can be directly applied. Fortunately, we find that a key feature of SLID is the active engagement of the liquidity pool owner, which creates the illusion of an active and reliable pool while gradually draining profits. Based on this, we generalize (via an empricial study) this single instance into a broader scam pattern that fully captures their distinct features.
5.1 Identification of SLID Pools
Given the lack of prior research on SLID scams, we used rug pulls as an intermediary to collect and analyze data from SLID pools in this pilot study. Our hypothesis was largely based on the similarities between SLID and rug pull scams, particularly in their initialization patterns and profit-taking actions, such as selling and withdrawing liquidity. We also hypothesize that some pools reported as rug pulls may actually represent potential SLID scams.
We first gathered rug pull reports dating back to 2022. Reliable sources for these reports included well-known crypto scam platforms such as Chainabuse [38], De.Fi [39], Chainanalysis [40], and scam warning accounts on X (formerly Twitter), such as ZachXBT [41] and Pocket Universe [42]. Our data collection yields a total of 921 reported rug pulls.
We then designed additional filtering mechanisms to enhance data validity and identify pools displaying SLID characteristics. The reported liquidity pools underwent a filtering process through four conditional layers in the following order:
  • $\bullet$
    Owner profit-checking layer. This layer computes the realized profit of the pool owner based on trading history data and excludes liquidity pools with negative realized profit. Since profit is a vital signal of the scam, this layer ensures the accuracy of the reports by filtering out the falsified reports where the token itself was not fraudulent but reported as a scam. As a result, 23 tokens were excluded from the dataset, specifically those associated with liquidity pools where the owner showed no profit. To be noted, this filtering process may also inadvertently exclude instances of failed scams or pools where the owner benefits through addresses not directly associated with the creator.
  • $\bullet$
    Honeypot layer. As the honeypot is relatively easy to detect during the pool deployment phase, we filtered out instances that are honeypot but reported as rug pull in our collecting process by utilizing GoPlus [43] and De.Fi [39] services. In this layer, we excluded 139 honeypot liquidity pools, which goes against the SLID predetermined characteristic of not having the token’s smart contract interference.
  • $\bullet$
    Mislabelled pools filtering layer. This layer processes the remaining liquidity pool instances and cross-references them with the rug pull detection methods proposed in previous studies [24, 37]. Out of these, 665 instances were confirmed as rug pull and filtered out, while 94 liquidity pools were identified as not being rug pull despite being reported as such on the online resources (claimed in §1).
  • $\bullet$
    Owner action layer. Amongst the 94 remaining instances, we recognize that they may contain various other frauds. To only select scams that align with our motivation example and the study purpose on SLID scams, we selected liquidity pools without the initial LP burned (which pool exists with the risk that the owner has all right to drain the pool at any time after getting investment, discussed at §2) and having at least three DEX activities committed by the owners, including deposit, withdraw, buy, or sell. This filter ensures that the remaining liquidity pool is aligned with our SLID predefined characteristics (§3) of having multiple activities executed by the pool owner. As a result, we excluded 23 liquidity pools and identified 71 pools that passed all four layers as our SLID initial ground truth for RQ1.
5.2 Features Finding of SLID Pools
We then conducted a comprehensive analysis of the filtered 71 liquidity pools and identified three common characteristics (C1, C2, and C3) of SLID pools. For comparison, our analysis included rug pull tokens that had been filtered out by the mislabeled pools filtering layer5.1). We also adopted rug pulls as the benchmark due to the similarities between the two scam types and obtain the distinct features that set SLID scams apart. Our analysis compared LPs from these two scam types across the following three aspects.
Figure 3: 30-day liveliness of rug pull and SLID LPs.
Post-month liveliness analysis.  One of the key metrics that can accurately distinguish SLID scams from rug pulls is post-month liveliness behavior. Cernera et al. [37] found that nearly 60% of the rug pull liquidity pools ended up “inactive” after one day of LP deployment. They further explored LP lifetimes for both scam types over an extended period of one month (cf. Figure 3), providing insights related to SLID scams as well. They revealed that 72.85% of SLID LPs remain active after the first month of deployment, whereas only 10.8% of rug pull LPs persist beyond this threshold.
To further investigate the differences in liveliness between the two scam groups, we compare the average changes in user activity for liquidity pools that remain “active” after the first month. On the first day of the second month, trading activities for the remain-active SLID liquidity pools were stable, showing an average reduction of 38.16% compared to the pool’s previous peak day with the highest recorded activity count. In contrast, the remaining rug pull liquidity pools experienced a dramatic decrease in recorded activities on the first day of the second month, averaging a 95.82% reduction. While other studies on cryptocurrency frauds (e.g., [37] on rug pull, [44] on Ponzi scheme, [45] on NFT rug pull) and detection tools tend to rely on the token’s liveliness as a vital factor for identifying scams, this metric appears to be ineffective in the SLID case due to its strong survival trait.
Our finding in SLID pool’s liveliness led to the following conclusion of SLID characteristic:
  • C1.
    The SLID liquidity pools are more adept at remaining undetected due to their prolonged liveliness. SLID scam pools remain active over a longer duration compared to rug pull despite using similar pool initiation and profit-taking mechanisms, making it nearly impossible to detect based solely on liveliness status.
Scammer activity analysis.  We also analyzed the scammer’s activity patterns in all targeted liquidity pools after the first month to learn the owner’s financial behaviors and motivations behind both fraud types. Given that the scammer is also the owner of these LPs, after victims buy paired tokens and send base tokens in the pool (pool value thus increased), scammers behind both rug pulls and SLID scams typically utilize two mechanisms for profit extraction: $(i)$ withdrawing liquidity from the pool and $(ii)$ selling the inflated token amount they hold externally to drain the liquidity pool.
Rug pull scammers prioritize immediate profit-taking, with 99.7% (663/665) of liquidity pools being entirely drained within the first day, and 100% drained within the first three days. The pool owners, who are also scammers, perform 75.38 DEX activities to their liquidity pool and 15.79 profit-taking actions on average during the first month. Among these profit-taking actions, each pool includes a single, large-amount profit-taking transaction (either selling or withdrawing) that constitutes over 95% of the total pool value. On average, those scammers extract 99.33% of the liquidity pool’s value after the first month, leaving only 0.67% as unrealized profit.
Scammers in SLID pools use contrasting strategies, operating more gradually and performing multiple profit-taking activities with smaller proportions. We observed that none of the SLID pools were completely drained within the first three days after deployment, in stark contrast to the 100% of all pools that were totally drained in rug pull scams during the same period. On average, SLID scammers executed 996.59 total orders during the pool’s first month, which was over 10 times more than the rug pull pool owners. SLID owners had a mean of 423 profit-taking actions, approximately 26 times more frequent than rug pull scammers. Additionally, SLID scammers control the proportion of profit-taking orders in a smaller range. The amount drained value from their profit-taking varies from a minimum of 7.39% to a maximum of 42.93% of the liquidity pool’s total value on average. This controlled proportion enables SLID scammers to evade detection by rug pull detectors, which typically focus on identifying pools with significant exit orders initiated by the owner.
To better demonstrate the differences in scammer actions between two scam types, we further analyzed (Figure 4) the impact of profit-taking activities by scammers in the MIND token [46] (rug pull liquidity pool) and the EMILY token [47] (confirmed inflated mint for rug pull by GoPlus [26] and found in our ground truth of SLID in §5.1) during their first three days. While a single profit-draining action by the MIND token owner completely emptied the liquidity pool and reduced the token’s price to zero, the EMILY token owner conducted multiple small-proportion exits that do not heavily affect the token’s price, thereby retaining future buyers’ confidence. Although both scam types initiate fraud similarly by creating liquidity pools and profiting through draining them via sell or withdraw orders, we observed that variations in execution duration and order proportions lead to distinct strategies for each scam type. From our analysis, we can conclude that:
  • C2.
    Although SLID scams employ a mechanism similar to that of rug pull, they pursue an opposite strategy characterized by a gradual approach, avoiding recognition. Scammers prioritize multiple small-proportion and slow-paced profit extraction activities, making it more difficult to detect, either through observation or by using detectors.
Figure 4: Scammer’s profit-taking activities and its effect on the token price of EMILY (SLID) and MIND (rug pull) liquidity pool in the first three days.
Profit and escalating potential analysis.  Based on the previous observation, the SLID strategy is expected to yield a better profit compared to rug pull by allowing scammers to conceal their activities while gradually attracting more victims. We validate this speculation via a profit-changing analysis between the two scam types. Technically, we focus on the realized profit and the unrealized profit of the LP owner.
We found that the realized profits (the actual profits sold or withdrawn by scammers from the pool) significantly differ between the two scam types. SLID scams demonstrate a superior profit margin: in the first month, while rug pull scammers achieve an average realized profit of 7.5 times their investment through a single, short-term draining action, SLID liquidity pools attain a massively higher profit margin of 10.3 times their investment through multiple small draining activities.
As for unrealized profits, SLID scames also perform a notable gap in profit advantage. Rug pull scammers have nearly zero future profit while completely emptying their liquidity pools, retaining only an average of 0.67% of pool value. In contrast, the SLID pool owner’s drainable assets in the liquidity pool after the first month retain, on average, 156% worth of their initial investment. This allows the SLID scammer, after the first month, to either fully drain the unrealized profit, resulting in an additional 1.56 times profit on top of the realized profit (already taken), or to refrain from withdrawing and let the pool continue to grow.
This observation indicates that by performing a vast number of small-proportion profit-taking orders (§3), SLID scammers manage to attract more victims, resulting in a larger netflow of funds poured into the pool. As a result, although they only take parts of their possible profit in the first month, these scammers already achieve better profit margins than rug pull scammers. Furthermore, a significant portion of assets that remain in the SLID pools retain substantial profit-scaling potential as the pool continues to grow, entrapping more future buyers. Therefore, we conclude that:
  • C3
    SLID scams offer a higher profit margin and sustainability, providing both significant mid-term profits and future profit-scaling potential if undetected. Compared to rug pull, SLID scammers can achieve a substantial growth rate in pool value from their initial investment, making it a more lucrative and enduring scam approach.
@@ 5.2 Features Finding of SLID Pools
6 Rule-based Detection for SLID Scams (RQ2)
The filters from RQ1 were explicitly designed to separate SLID from rug pull LPs, however, they may perform ineffectively for identifying SLID scams from the population containing various types of liquidity pools. We thus propose a rule-based heuristic designed to effectively detect SLID LPs among the global dataset of Ethereum pools.
6.1 Rule-based Heuristic for SLID Detection
Overview. Our heuristic consists of three validators: Honeypot Validator, profit Validator, and Owner Activity Validator. Each validator targets different characteristics from the extracted data of LPs (cf. RQ1 findings). If one LP is flagged by all three validators (i.e., exhibiting all observed traits), we add it to the initial ground truth dataset as a SLID scam.
Honeypot validator.  This component assesses whether tokens are honeypots, as these are inconsistent with the typical characteristics of an SLID scam. A successful SLID scam often depends on a gradual increase in victim participation and buying volume, fueled by the belief in the token’s legitimacy (see C2). Therefore, it’s crucial for an SLID liquidity pool to display standard initial DEX activities, such as buying and selling, without indications of a honeypot.
Our lightweight Honeypot Validator employs rule-based criteria derived from prior research of [25, 48]. Instead of simply replicating exact detection methods, we apply suggested criteria to specific features extracted from the token’s smart contract and set an upper threshold to identify potential honeypot signals.
In particular, we extracted the following features from the paired token’s smart contract:“buy_tax”, “sell_tax”, “cannot_buy”, “cannot_sell_all”, “slippage_modifiable”, “transfer_pausable”, “personal_slippage_modifiable”, and “trading_cooldown”. These features were evaluated based on specific criteria; each indicates different scenarios in which the pool’s owner can illegally interfere with the user’s activity in the pool for gaining profit. An upper threshold was defined to identify potential honeypots.
Here, we provide an intuitive example: a liquidity pool is flagged as a honeypot if the token’s “buy tax” exceeds 50%, based on tax-related criteria from earlier research [25].
Profit validator.  Profit-related activities reveal the SLID characteristics (§4) from two critical insights (short for I): (i) I-1: the scam likelihood by evaluating realized profit, an essential element of all successful scams; (ii) I-2: the scam likelihood by evaluating the patterns of unrealized profits generated. Thus, we set both realized profit and unrealized profit as our metrics for analyzing the owners’ profit behaviors. Those values can be obtained through the collected DEX orders of liquidity pools (more in §7).
The pool owner’s realized profit can be computed by taking the difference between the total value of their investment in the pool (in USD) and the total value that has been withdrawn over the pool’s lifetime under the base token (detailed equation in Appendix C). We flag a liquidity pool where owners have a positive total realized profit from their activities. This reflects the scam likelihood in terms of profit generation, per I-1. Although not all projects that generate realized profits are scams, our observations in §4 show that all reported scams exhibit a positive realized profit. This indicator helps the heuristic identify and exclude projects with no profit.
The unrealized profit of the pool’s owner is another critic metric, especially when the SLID LP shows a significant growth in unrealized profit after the first month of deployment. The first-month unrealized profit represents the portion of profit (in the base tokens) belonging to the owner from the liquidity pool’s value after its initial month of deployment (detailed computing equation in Appendix C). In §4, we applied this unrealized profit metric for comparisons between SLID and rug pull scams, revealing a significant disparity between the average first-month unrealized profit ratios generated by the two groups (0.67% of the initial investment for rug pull and 156% of the initial investment for SLID pools). However, we set the threshold for the unrealized profit in this heuristic to 0 due to the uncertainty stemming from our current small initial ground truth dataset from RQ1, with the lowest computed unrealized profit for a SLID instance being $1.3e^{-14}$. This threshold addressed our I-2 of the scam likelihood pattern in generating unrealized profit.
We acknowledge that limiting the capturing range of unrealized profit to only within the first month might exclude potential SLID cases where the owner’s pool share develops afterward. Thus, a soft heuristic threshold of first-month positive unrealized profit allows a return with a lower bound on the ground truth dataset that only collects relevant cases with strong similarity to SLID discovered characteristics.
Owner activity validator.  SLID scams prioritize a gradual approach by multiple small-proportion profit-taking actions over an extended period (§5). We accorndingly built Owner Activity Validator in the following steps:
  • $\bullet$
    excludes the liquidity pools where the pool owners faithfully burn their initial (LP-) tokens after deploying the pool;
  • $\bullet$
    continue to detect the pools where their owners still getting all control over the initial (LP-) tokens;
  • $\bullet$
    observe/analyse the owner’s activities in exploiting profits.
We measure their activities through two metrics: (i) Profit-taking order count, denoted by $c$, accounts for the total number of DEX orders involved in profit extraction; and (ii) order impact, denoted by $I$, indicates the ratio of the $z^{th}$ profit-taking order size of value compared to the pool’s total value before the activity is executed. We define their thresholds as:
$c\geq{T_{count}},\,\text{where}\,I_{z}=\frac{y_{z}}{{x}_{t-1}}<T_{impact},\,\forall\ z\in\{1,2,\dots,c\}.$ (1)
where $T_{count}$ and $T_{impact}$ are the thresholds (§5) of owner’s profit-taking orders counts and maximum impact on the pool.
We select $T_{count}=5$ as the lower threshold for the profit-taking order count from the owner for their pool to be flagged as SLID. This choice is supported by our observations in §5: the SLID owner with the least profit-taking activities engaged in 5 profit-draining orders. Additionally, $I_{z}$ is computed as the order size, $y$, of the $z^{th}$ profit-taking order from the owner divided by the previous pool size ${x}_{t-1}$, where $t$ in this equation represents the order count of profit-taking order $z$.
To reflect the characteristic of small-proportion orders, we set the impact threshold of all owner’s profit-taking orders to be less than 95% ($T_{impact}<0.95$) of the total pool size. Our analysis in §5 revealed that 95% was the lowest profit-taking impact recorded for a rug pull instance. It also ensures we capture a broader range of scams, including potential SLID scams that have been unnoticed by rug pull detection heuristics. In comparison, for rug pull detection heuristics in [24, 37], only pools with a 0.99 profit-taking impact are flagged.
Additionally, we decide not to incorporate adaptivity for all numerical thresholds stems from the lack of ground truth in this new exploration of SLID . We focus on confirmed scams, including mislabeled rug pulls. We acknowledge that in real-world systems, attackers may adapt to the threshold and evade detection, but our model still lays a foundation.

### v3

@@ II Technical Warm-Ups
Honeypot. Another typical scam is Honeypot [25]. The scammer maliciously designs the token’s smart contract and allows users to easily purchase tokens but hardly sells them. The contract may include restrictions such as preventing sell transactions, allowing withdrawals only for whitelisted addresses, or imposing exorbitant transaction fees that make it impossible to retrieve funds. The scammer then exploits LPs through sell/withdraw actions, similar to the mechanics of rug pull.
We defer more related work to Appendix A-A.
@@
Without implementing an ethical safeguard (i.e., burning initial LPTs), liquidity providers, including those involved in rug pulls, honeypots, and SLID scams, confront no financial risk (proven in Appendix A-B) for engaging in fraudulent activities.
III Motivating Examples
III-A Case Study of Initial SLID Liquidity Pool
We provide a typical case study. We discuss the behavior of Token $\hat{A}$ (here, our collected data, including token names, addresses, and LP details, are anonymized for ethical reasons) with token address “0x4e84…a17c8”, created in August 2021, which serves as a motivating example for our study. This token was launched with a massive marketing campaign highlighting its high potential for generating profits for investors, and later got reported by the public (details in §IV-A) for the owner’s fraudulent selling activities.
We identified the LP of $\hat{A}$ as a SLID scam due to the deceptive actions of its owners. They slowly exploited the LP through a mixed sequence of token buying and selling on Uniswap while maintaining a stable price and trading volume during the initial months after its launch. The gradual profit-taking persisted for an extended period before the token owner executed a significant withdraw a year later. We briefly describe the sequence of actions taken by the owners in this cryptocurrency’s SLID scam.
➊ Attracting victims. To rapidly attract users, $\hat{A}$’s advertising campaign targeted buyers by promising substantial token profits if purchased early. The owner engaged the community by promoting the token via their X account (formerly Twitter). They attracted 78.5K followers and collaborated with 20+ well-known crypto influencers. The campaign drew 5,326 investors to purchase $\hat{A}$ on Uniswap within the first 70 days from the token’s LP deployment.
Figure 1: SLID workflow: ① create a liquidity pool; ② add tokens (base & paired); avoid burning initial shares; ④ marketing campaign; ⑤ invest funds and buy tokens; ⑥ normal activities (buy orders); slowly withdraw/sell funds; ⑧ liquidity drop; money loss.
➋ Maintaining the LP and slowly taking profit.  The LP was deployed on Aug. 4th, 2021, with an initial owner liquidity deposit of \$19K. However, the LP for $\hat{A}$ was deployed without burning the owner’s initial tokens. This creates a persistent risk, as the owner of $\hat{A}$ can drain the LP and seize all funds invested by users to purchase $\hat{A}$ at any time. This is precisely what occurs in rug pull and honeypot scenarios (cf. §II).
We observed $\hat{A}$’s owner DEX activities (Figure 2) over 70 days after deployment, including price and trading volume. During this period, the owner was engaged in a sequence of mixed activities, including buying and selling. Although these actions seem legitimate, our evaluation reveals that the selling amount consistently exceeded the buying amount (in paired tokens). Specifically, the total selling volume was ten times larger, with the owner purchasing only $2.38\times 10^{15}$ tokens while selling $3.21\times 10^{16}$ tokens of $\hat{A}$ in total. This indicates that the owner possesses a significant number of tokens and attempts to incorporate this inflated amount into their selling orders. $\hat{A}$’s owner executed 139 buy and 136 sell actions, but with a double volume gap that was worth \$783K and \$1.4M, respectively. This resulted in a huge profit for the LP owner of \$707K — double the initial investment. In contrast, 5,326 investors committed with 10,374 buys, 3,898 deposits, and 7,322 sells, but only 2 withdraw activities, with a total inflow of \$1.35M (in base token) to the LP. The mix of buy and sell orders created the appearance of legitimacy in the token’s transaction history, helping to avoid the token price crash that would occur if they sold all at once (i.e., 1-day rug pull [39]).
➌ Repeating the scam. While maintaining the appearance of token legitimacy, $\hat{A}$’s owner continued to sell only a small portion of the tokens compared to the overall LP value. The smallest and largest sell order values account for only 1% and 16% of the LP’s value. By carefully mixing buy and sell orders and controlling the selling amounts, the built-in scam detectors on DEXs are unable to identify this token as a scam. As of July 13, 2024, the scammer still maintains \$91K worth of liquidity in the LP and has remained inactive since their last recorded sell order on July 31, 2022.
Thus, despite the token’s stable behavior, a series of buy and sell actions (with smaller amounts) occurred at a slow pace, gradually adding an inflated amount of paired tokens (§V-A) to each sell order from the LP owner’s account. This scam strategy can easily avoid investor attention while enabling tokens to remain active and attract additional investments over an extended period. The transaction history of $\hat{A}$’s LP on Uniswap in the analyzed period is detailed in [40].
III-B More LPs Exhibiting SLID Patterns
In addition to the motivation example, we present more LPs, i.e., $\hat{B}$ and $\hat{C}$, exhibiting similar operational patterns. All those projects receive community public reports (§IV-A).
LP of token $\hat{B}$. LP $\hat{B}$ (with paired token address “0x59…e9CD”) followed a suspiciously similar lifecycle, beginning with the initial deployment of an LP on 12 Feb 2024 without burning LP tokens, granting persistent unilateral control to its creators. Our analysis of transaction records revealed that the owner obtained a huge amount of inflated paired tokens minted outside and continuously sold them into the LP from 12 February to 17 April in a moderate portion. The owner incrementally extracted investor funds of \$2660.9 after the first 2 months of the LP’s deployment.
Notably, up to today, token $\hat{B}$’s trading activity in the same LP showed frequent but minor liquidity withdrawals interspersed with small buy and sell orders, similar to $\hat{A}$. The only difference is that $\hat{B}$ does not run a large marketing campaign. Despite attracting fewer buyers, the owner still generated illicit profits.
LP of token $\hat{C}$. LP $\hat{C}$ (token address “0xbA…9ccE”) also exhibited trading patterns indicative of SLID behaviors. The LP was deployed on 7 April 2024. The owner maintained control without burning LP tokens, possessed inflated tokens externally, and gradually sold them with 38 sell orders from April 9 to April 26, 2024. Meanwhile, LP $\hat{C}$ displayed frequent, small-scale liquidity withdrawals combined with sporadic minor transactions. The LP only employed minimal promotional activity from X’s influencer. Unlike other examples, this scam resulted in a realized profit of \$200 for the owner. The realized profit appeared to be smaller compared to other scams during the same period. $\hat{C}$ transaction history still illustrates the core SLID behavior (more details in §V-A).
All the transactions involved in the motivation examples analysis of $\hat{A}$, $\hat{B}$ and $\hat{C}$ are provided in [40].